OpenStudy (anonymous):

sum (k^2-4),k=0..14 the answer is 955 just give me the steps please

OpenStudy (anonymous):

14 Σ (k^2-4) k=0

OpenStudy (anonymous):

\[\sum _{k=0}^n k^2=\frac{1}{6} n (n+1) (2 n+1)\\ \sum _{k=0}^{14} k^2=\frac{1}{6} n (n+1) (2 n+1)=1015\\ 1015 - 15 (4)= 955 \]

OpenStudy (kinggeorge):

\[\sum_{k=0}^{14} (k^2-4)=\sum_{k=0}^{14} k^2 - \sum_{k=0}^{14} 4\]\[=\sum_{k=0}^{14}k^2 -15(4)\]\[={14(14+1)(2(14)+1) \over 6} -60=1015-60=955\]